How many moles of oxygen are produced when 756 grams of lead 11 oxide are formed

Find the mass of nitrogen dioxide that forms with 0.58 moles of oxygen
1
Expert's answer
2020-04-13T02:29:03-0400

1) 2 Pb + O2 = 2 PbO

the amount of PbO is 756 g / 223.2 g/mole = 3.38 mole, it is twice of moles of O2, i.e.

the amount of oxygen is 3.38/2 = 1.69 moles

2) N2 + 2 O2 = 2 NO2

0.58 moles of oxygen correspond to 0.58 moles of nitrogen dioxide,

the mass of nitrogen dioxide is 0.58 mole * 46 g/mole = 26.68 g